# Dynamic Indexing Through Learned Indices with Worst-case Guarantees

### v1

@@ 6 From an $\varepsilon$-cover to an indexing structure
A learned index $h$ does not immediately support indexing and range queries. We obtain an indexing structure by combining $h$ with a hash map $H$. Combining learned models with hash maps is not new [20, 28, 30] and this technique has even been applied to learned indexing [19]. The core idea is to store $S$ in an unordered vector $A$ and maintain a Hash map $H:\mathbb{Z}\mapsto[n]$. Given some $q\in\mathcal{U}$, the learned function then produces a value $v$ such that $A[h(v)]$ is ‘close’ to $q$. In our setting, it is compelling to create $H$ such that $A[H(h(q))]$ is (approximately) the predecessor of $q$. However, dynamically, this approach fails for the same reason that storing each $s\in S$ at $A[\textnormal{{rank}}(s)]$ fails. Since the ranks of elements in $S$ are constantly changing, we build a hash map using the part of $S$ that remain constant: the values.
@@
In Appendix B, we define a data structure independent of the learned index $h$ (Figure 5 (a) + (b)). A page $p$ is an integer with a vector that stores all $s\in S$ where $\lfloor\frac{s}{\varepsilon}\rfloor=p$, in order. We store all non-empty pages $P$ in an unordered vector $A$. We maintain a hash map $H:P\rightarrow[|A|]$, where $A[H(p)]$ contains the page $p$. We additionally maintain a doubly linked list over all pages in $P$, arranged in sorted order.
@@
We restrict our $\varepsilon$-cover $f$ to a vertical $\varepsilon$-cover. It is a $y$-monotone collection of line segments such that for all points $p\in F_{S}$, a vertical line segment of height $2\varepsilon$ centred at $p$ intersects a segment in $f$. We compute $f$ oblivious of our paging structure.
Given $q\in\mathcal{U}$, we project $q$ onto $f$ (Figure 5 (d)). From there, we project to the $x$-axis, floor the value, and project back to $f$. We prove that the resulting $y$-value corresponds to the page $p$ containing predecessor($q$). This way, we answer predecessor using $O(\varepsilon+\log|f|)$ time. We prove in Theorem B.11 that gives our desired update and query times. Note that our approach increases $|h|$ as we obtain the cover from a more restricted class of $\varepsilon$-covers.
@@ 7 Experiments
Our implementation is written in C++ (anonymous.4open.science/r/DynamicLearnedIndex). We compare to the C++ implementation in [13], which uses a PGM index under the logarithmic method. The experiments were conducted on a machine with a 4.2GHz AMD Ryzen 7 7800X3D and 128GB memory. Our test bench is available at anonymous.4open.science/r/LearnedIndexBench. It can replicate experiments, generate synthetic data, and produce plots. As input we consider two synthetic data sets and two real world data sets, each category containing one set of geometric nature and one of random nature. The data sets vary in size and have differing patterns and structure, showcasing performance in various settings. Each set consists of unique 8 byte integers in randomly shuffled order.
  • LINES is a synthetic data set of 5M integers that, in rank space, produce $m$ lines of exponentially increasing slope. This set models the ideal scenario for a PGM index.
@@
We compare the quality of the learned indices based on space consumption and the complexity of $h$, in a dynamic setting. We use the same choice of $\varepsilon=64$ as in [13] across our experiments. For performance, we focus on range queries. We measure their time per operation and space usage in a dynamic setting. Our summary statistics consist of the number of lines stored in the learned index $h$ throughout construction, and the space consumption of the search structure over $h$. We note that logarithmic PGM is by default equipped with an optimisation that avoids building a PGM for data below a certain size. In this case, it instead only uses an underlying sorted array without additional search structure. In order to properly compare the performances, this optimisation has been disabled.
Following dynamic precedent.
@@
For the dynamic scenarios, we first follow the precedence set by prior papers [13, 16]. These first construct, insertion-only, the indexing structure. They then perform a batch of $10$M operations. These batches consist of insertions, deletions, and range queries over ranges such that the output contains approximately $\frac{\sqrt{n}}{10}$ elements. We vary the ratio of queries in a batch and evenly split the remaining operations between insertions and deletions. The batch is shuffled before processing. We deviate from this procedure for LINES. The goal of LINES is to illustrate how these algorithms compare under (near)-perfect structure. Therefore, we keep the structure of this data set largely intact and restrict the batch of updates to lie on only one of the line segments. We limit the number of operations in a batch to $1M$ to align the relative sizes between workloads and the data set.
Adversarial data.
@@
Counting segments.
Figure 6 shows the number of line segments stored by the dynamic and logarithmic PGM respectively we observe two very distinct patterns of behaviour between the geometric and non-geometric data. For the geometric data sets, we find that the dynamic and logarithmic PGMs perform similarly up to a certain point, at which the performance of the logarithmic PGM deteriorates. The logarithmic partitioning scheme is easily visible from the jagged nature of the measurements. In addition, because data remains scattered across $O(\log n)$ buckets, the logarithmic method fails to realise the optimal line count for the synthetic LINES data by a considerable margin. For the LONGITUDE data the logarithmic PGM maintains approximately fifty percent more lines as the dynamic PGM.
For the UNIF and BOOKS data, which lack inherent geometric structure, we observe very similar trends between the dynamic and logarithmic PGMs, with the logarithmic PGM maintaining around thirty percent fewer lines as the dynamic PGM. It appears that on unstructured data, a random partition of the data makes the result more structured.
Figure 6: The number of lines stored in the learned indices, insertion-only. The top measurements are from geometrically structured data. The left measurements are from synthetic data.
Operational performance. For the dynamic scenario, shown in Figure 7, we measure the total running time of processing batches of $10$M operations. The $x$-axis shows the ratio between queries and updates within the batch. The rightmost column is our adversarial case.
In these results, there are also distinguishing patterns between geometric and non-geometric data. For the geometric data, we pay a much higher cost for updates to the structure relative to the cost of queries. This is easily seen on the LONGITUDE data, where there is a steep decrease in running time as the query ratio increases. The cost of updates stems from having to test if lines containing many points can potentially be merged. While the update time is $O(\log^{2}n)$, the memory access pattern is very bad. It incurs cache misses as tree structures have to be split, joined, and the internal structure is restored. This is also the reason why we do not see this behaviour on the LINES data, as the memory overhead is small enough that the latency is hidden.
For the random data, the performance instead increases with the query ratio. As the segments here contain fewer elements, the impact of the previously mentioned tree structure operations is limited. However, for range queries, the dynamic PGM here is heavily penalized by the unordered nature of the underlying array. When reporting a range, the dynamic PGM potentially occurs a cache miss for every element reported. The low memory overhead, and contiguous storage, of the logarithmic PGM puts it far ahead on performance in the random setting. Notably, as the sizes of data sets increase, the logarithmic PGM is also affected by memory access, which is most evident on the LONGITUDE data.
Adversarial data.
Across all data sets, the effects of the tombstoning strategy of the logarithmic PGM become critically apparent in our adversarial scenario (denoted by ADV in Figure 7). The dynamic PGM is an order of magnitude faster on all data sets. This occurs because a range query for the logarithmic PGM iterates over all deleted values between the query parameters $(q,t)$, subtracts their tombstones, and then outputs the desired range. Our approach is instead output-sensitive.
@@
Figure 7: Time per operation in dynamic scenario with 10M operations with varying query ratios. For the LINES data set the batch size is $1$M, and updates are restricted to points on a single line.
Summary statistics.
We also compiled summary statistics in Table 2 to give insight into other metrics of comparison. For the geometric data we observe that we pay a high price of updates, with construction times two orders of magnitude greater than that of the logarithmic PGM. On the positive side, we maintain roughly half as many segments at the maximum. For random data. the construction time is approximately one order of magnitude greater, with an increase of around thirty percent in the number of segments maintained. In terms of space consumption, the dynamic PGM uses nearly 7 times the memory of the logarithmic PGM, on geometric data, and close to 8.5 times on random data. For the adversarial scenario, the memory consumption is comparable or slightly better for the dynamic PGM.
Index data set Construction Max no. of lines Space Adv. space
Logar. PGM LINES $2$S $204$ $80$MB $26$MB
UNIF $22$S $3.393$ $800$MB $263$MB
BOOKS $48$S $5.978$ $1.600$MB $527$MB
LONGITUDE $132$S $8.227$ $1.663$MB $831$MB
Dynamic PGM LINES $415$S $93$ $520$MB $17$MB
UNIF $199$S $4.238$ $6.795$MB $2.147$MB
BOOKS $594$S $7.158$ $13.613$MB $537$MB
LONGITUDE $13.363$S $4.319$ $11.076$MB $339$MB
Table 2: Summary statistics of the learned indices for a variety of data sets. Adversarial space refers to the memory consumption after processing the adversarial dynamic scenario. Construction time and adversarial space consumption are measured for the indexing structures.
@@ 8 Conclusion
In this paper, we study learned indices under a geometric lens. The rank-space of a data set is a two-dimensional Euclidean space, and we use tools from two-dimensional Euclidean convex hull computation to improve on algorithms that are popular for dynamically maintaining rank-space information. Our original aim was to answer two questions:
@@
“Can a learned indexing structure be dynamically maintained with worst-case guarantees?”
We provide new algorithms for dynamically maintaining a learned index and an indexing structure. Here, we utilise existing techniques from computational geometry in this new setting. While robustness is usually the crux of computational geometry implementations, we observe that it is a challenge to implement convex hull intersection testing in a robust manner. We give a new domain-specific implementation for convex hull intersection testing and prove its correctness. We extend this algorithm to compute a separating line between the hulls in the negative case.
Our analyses and implementations provide insights into robust convex hull intersection testing and illustrate how dynamic convex hull data structures can be applied in practice. Our empirical analyses indicate that our learned index $f$ uses considerably fewer line segments on geometric data, and a comparable amount of line segments on arbitrary data.
We combine our learned index with techniques from randomised data structures, i.e. hashing, and with a doubly linked list data structure. We obtain a dynamic indexing structure with output-sensitive worst-case guarantees, even when the data set is subject to deletions. Due to the nature of our memory consumption, and the access patterns in the tree structures, our approach is often slower in practice. The fact that the logarithmic PGM has contiguous memory access offsets the cost of having to query $O(\log n)$ data structures. We observe that for geometric data, our running times are competitive and become increasingly favourable as the query-to-update ratio increases. Our data structure is considerably more efficient across all data sets in an adversarial scenario in which there are many deletions (followed by a batch of queries). While this scenario may perhaps not be realistic, it does illustrate the worst-case guarantees of our implementation. We consider it an interesting open problem to see whether our tree-based solution can be combined with array-based implementations to improve the memory access.

### v2

@@ 4 Robustness
5 Dynamically maintaining a learned index
We dynamically maintain a learned index $h_{\varepsilon}$ of $S$ by maintaining an $\varepsilon$-cover $f$ of $S$. We guarantee that there exists no $\varepsilon$-cover $f^{\prime}$ of $S$ with $|f|>\frac{3}{2}|f^{\prime}|$. By Observation 2, we obtain a learned index $h_{\varepsilon}$. By Observation 2, there exists no PGM index $h_{\varepsilon}$ where $|f|>\frac{3}{2}|h_{\varepsilon}|$.
Theorem 5.1.
Proof 5.2.
The proof is illustrated by Figure 4. For any $s,t\in\mathbb{Z}$ with $s\leq t$, we say that $S[s,t]$ is blocked if there exists no $\varepsilon$-cover of $S[s,t]$ of size $1$. We maintain an $\varepsilon$-cover $f$ where for all consecutive intervals $[a,b],[c,d]\in\Lambda(f)$, $S[a,d]$ is blocked. Thereby, $|f|\leq\frac{3}{2}|f^{\prime}|$ for any $\varepsilon$-cover $f^{\prime}$ of $S$ (we give a proof of this fact in Appendix C).
[…]
Let $([w,x],[y,z],[a,s],[s,s],[s,b],[c,d],[e,f])$ be consecutive intervals in $\Lambda(f)$ and denote $I=([y,z],[a,s],[s,s],[s,b],[c,d])$ (see Figure 4 (c) ). For each $(s,t)\in I$, we have access to $T(S[s,t])$. For any consecutive pair $([s,t],[q,r])$ in $I$, we may join the trees $T(S[s,t])$ and $T(S[q,r])$ in $O(\log^{2}n)$ time to obtain $T([s,r])$. We invoke $T([s,r])$.get_hull() and apply Theorem 3.4 to test in $O(\log^{2}n)$ total time whether $S[s,r]$ is blocked. If it is not, we replace $[s,t]$ and $[q,r]$ by $[s,r]$. Otherwise, we keep $T(S[s,r])$ and a complexity-1 $\varepsilon$-cover of $S[s,r]$.
@@ 6 From an $\varepsilon$-cover to an indexing structure
A learned index $h_{\varepsilon}$ does not immediately support indexing and range queries. We obtain an indexing structure by combining $h_{\varepsilon}$ with a hash map $H$. Combining learned models with hash maps is not new [26, 37, 39] and this technique has even been applied to learned indexing [25]. The core idea is to store $S$ in an unordered vector $A$ and maintain a Hash map $H:\mathbb{Z}\mapsto[n]$. Given some $q\in\mathcal{U}$, the learned function then produces a value $v$ such that $A[h(v)]$ is ‘close’ to $q$. It is compelling to create $H$ such that $A[H(h(q))]$ is (approximately) the predecessor of $q$. However, dynamically, this approach fails for the same reason that storing each $s\in S$ at $A[\textnormal{{rank}}(s)]$ fails. Since the ranks of elements in $S$ are constantly changing, we build a hash map using the parts of $S$ that remain constant: the values.
@@
In Appendix B, we define a data structure independent of the learned index $h_{\varepsilon}$ (we illustrate our approach in Figure 5 (a) + (b)). A page $p$ is an integer with a vector that stores all $s\in S$ where $\lfloor\frac{s}{\varepsilon}\rfloor=p$, in order. We store all non-empty pages $P$ in an unordered vector $A$. We maintain a hash map $H:P\rightarrow[|A|]$, where $A[H(p)]$ contains the page $p$. We additionally maintain a doubly linked list over all pages in $P$, arranged in sorted order.
@@
We restrict our learned index $h_{\varepsilon}$ to a vertical $\varepsilon$-cover. I.e., $h_{\varepsilon}$ is a $y$-monotone collection of line segments such that for all points $p\in F_{S}$, a vertical line segment of height $2\varepsilon$ centred at $p$ intersects a segment in $h_{\varepsilon}$. We compute $h_{\varepsilon}$ oblivious of our paging structure.
Given $q\in\mathcal{U}$, we project $q$ onto $h_{\varepsilon}$ (Figure 5 (d)). We project to the $x$-axis, floor the value, and project back to $h_{\varepsilon}$. We prove that the resulting $y$-value corresponds to the page $p$ containing predecessor($q$). This way, we answer predecessor using $O(\varepsilon+\log|h_{\varepsilon}|)$ time.
As a result, we dynamically maintain a learned index $h_{\varepsilon}$ and a data structure that updates in $O(\varepsilon+\log^{2}n)$ and supports indexing queries in $O(\varepsilon+\log|h_{\varepsilon}|)$ expected time (Theorem B.11).
@@
Theorem 6.1.
For any $\varepsilon$, there exists a data structure to dynamically maintain a vertical $\varepsilon$-cover $F$ of a dynamic set of distinct integers in $O(\varepsilon+\log^{2}n)$ time. We guarantee that there exists no vertical $\varepsilon$-cover $F^{\prime}$ with $|F|>\frac{3}{2}|F^{\prime}|$. The data structure supports indexing queries in $O(\varepsilon+\log|F|)$ expected time and range queries in additional $O(k)$ time where $k$ is the output size.
@@ 7 Experiments
Our implementation is written in C++ and made publicly available [20]. We compare to the C++ implementation in [17], which uses a PGM index under the logarithmic method. The experiments were conducted on a machine with a 4.2GHz AMD Ryzen 7 7800X3D and 128GB memory. Our test bench is available [21], and can replicate experiments, generate synthetic data, and produce plots. As input we consider two synthetic data sets and two real world data sets. Three contain data of geometric nature, with one of random nature to align with precedent. Each set consists of unique 8 byte integers in randomly shuffled order. In Appendix D, we showcase additional experiments on other datasets.
  • LINES is a synthetic data set of 5M integers that, in rank space, produces 5 lines of exponentially increasing slope. This set models the ideal scenario for a PGM index.
@@
  • DRIFTER is a real world data set, containing roughly 1.7M steps of accumulated distance travelled by ocean drifters tracked through GPS [8].
@@
We compare the quality of the learned indices based the complexity of $h_{\varepsilon}$, in a dynamic setting. We use the same choice of $\varepsilon=64$ as in [17] across our experiments. For performance of the indexing structures, we measure their time per operation in a dynamic scenarios with a range of query to update ratios. We note that logarithmic PGM is by default equipped with an optimisation that avoids building a PGM for data below a certain size. In this case, it instead only uses an underlying sorted array without additional search structure. In order to properly compare the performances, this optimisation has been disabled.
7.1 The learned index complexity
Figure 6 presents the complexity of the learned indices, measured by the number of line segments maintained during random-order insertions. The behaviour differs between datasets. For the geometric Lines and Longitude datasets, the dynamic and logarithmic PGMs initially perform similarly. As the data grows, the performance of the logarithmic PGM degrades and its jagged progression reflects its logarithmic partitioning.
On the highly structured synthetic Lines data, the logarithmic method consistently retains more segments than necessary. It misses the optimal line count by a wide margin due to its fragmentation across $O(\log n)$ buckets. A similar pattern appears in the Longitude dataset, where the logarithmic PGM maintains roughly 50 percent more segments than our solution. We note that precisely on these structured data sets, the complexity of the learned indices is $o(n)$. I.e., precisely here one also expects improvements in query time.
Figure 6: The complexity of the learned indices throughout insertion-only construction. The top graphs represent geometrically structured data. The left graphs represent synthetic data.
For unstructured data such as Unif, both learned indices display similar asymptotic trends, with the logarithmic PGM using approximately 30 percent fewer segments. Here, the complexity appears to scale as $\Omega(n)$ for both methods, suggesting that learned indices offer few to no improvements in the absence of exploitable structure. Surprisingly, the Drifter data – despite its geometric origin – shows similar results to that of Unif. Again, complexity scales linearly, and the logarithmic method outperforms the dynamic one by around 30 percent. This implies that whatever latent geometric structure exists is insufficiently captured by the learned index under either strategy.
An interesting observation is that, on unstructured data, the imposed bucketing of the logarithmic PGM index can introduce a form of regularity that the model benefits from, essentially imposing artificial structure where none exists.
7.2 Running time comparisons
Recall that our indexing structure is composed of a learned index over a vertical $\varepsilon$-cover and a paging structure. We first examine the performance of both the learned index in its own, and then the performance of the full indexing structure.
@@ 7.2 Running time comparisons
For the dynamic scenarios, we first follow the precedence set by prior papers [17, 22]. These first construct, insertion-only, the indexing structure. They then perform a batch of $10$M operations These batches consist of insertions, deletions, and range queries over ranges such that the output contains approximately $\frac{\sqrt{n}}{10}$ elements. We deviate from the precedent by also deleting from the index before performing a batch of operations, to simulate a scenario in which the structure has existed and transformed prior to processing.
7.2.1 Maintaining a learned index
Figure 7 shows the cost of maintaining each learned index under dynamic operations. Our update procedure, which ensures worst-case $O(\log^{2}n)$ bounds, performs significantly worse than the amortised $O(\log n)$ updates offered by the logarithmic method – especially on larger datasets such as Longitude and Unif. On smaller datasets like Drifter, the gap narrows, but the logarithmic method generally remains preferable in these update-only scenarios.
Figure 7: Update times for maintaining the learned index dynamically.
7.2.2 Indexing data structures
We next assess the complete indexing structures under mixed workloads. Batches contain a tunable ratio of queries and updates, where updates are evenly split between insertions and deletions, and operations are randomly ordered. More experiments can be found in Appendix D. For the Lines dataset, updates are constrained to a single line segment to preserve its idealised structure. For smaller datasets, batch size is limited to 25 percent.
Adversarial workload (ADV).
@@ 7.2.2 Indexing data structures
Figure 8: Time per operation in dynamic scenario with varying query ratios. For the LINES data set updates are restricted to points on a single line. The ADV ratio denotes the adversarial case where all but a constant number of elements are deleted before the structure is queried.
Operational performance.
Figure 8 summarises the total processing time under varying query-to-update ratios. Structured datasets, particularly Longitude, reveal a trade-off between the fact that we have a lower-complexity learned index and our update time. On one hand, our updates are costly due to our line-merging tests which suffer from the poor cache behaviour from pointer-based trees, and complex rebalancing that is present in the state-of-the-art practical dynamic convex hull data structures. This behaviour is also reflected on our previous analysis of maintaining the learned index itself. On the other hand we see on geometrically structured data that, as the query ratio increases, performance improves. This trend is absent in the Lines data, where updates are constrained and the structure is small, mitigating cache penalties. Here, running time is largely driven by query processing, which increases as expected with query ratio.
In contrast, the random data from the Unif dataset shows improved performance with higher query ratios. In our data structure, having smaller segments reduce restructuring costs. Our dynamic implementation suffers greatly from cache inefficiency due to random access during range reporting. The logarithmic PGM, with its lower memory overhead and sequential layout, delivers significantly better performance – particularly as sizes increase.
For the Drifter data, one would expect a trend similar to that of the Unif data, based on the complexity of the $\varepsilon$-cover shown in Figure 6. However, there is little difference in performance for either PGM as the query ratio is varied. This is likely due to the small size of the data set, with both structures performing updates in fractions of microseconds on average. The logarithmic PGM does come out slightly on top, likely due to the machine friendly memory access pattern.
Adversarial impact.
Across all datasets, the logarithmic PGM’s tombstoning strategy becomes a bottleneck in the adversarial scenario. Its range queries must scan and subtract deleted values before outputting results. Our structure, by contrast, is output-sensitive and avoids such overhead. This illustrates that our data structure, whilst being generally less efficient that the logarithmic PGM, offers worst-case guarantees.
@@ 8 Conclusion
We studied dynamic learned indices through a geometric perspective. Following the work of Ferragina and Vinciguerra [17], we maintained a learned index $h_{\varepsilon}$ of a dynamic set $S$ as a piecewise-linear approximation—an $\varepsilon$-cover—of the rank-space point set $F_{S}$. We used techniques from computational geometry to answer the following question:
@@
We proposed a new approach to maintain a learned index based on dynamic convex hull data structures. We presented an $O(\log^{2}n)$ time algorithm to dynamically maintain a learned index $h_{\varepsilon}$. To obtain this algorithm, we showed an algorithm to compute a separating line between two non-intersecting convex hulls – an operation previously missing from the literature. The existing logarithmic PGM index has an amortised $O(\log n)$ time update algorithm which is more efficient in practice. Indeed, although close in theory, the memory access pattern associated with dynamically maintaining convex hulls incurs heavy penalties for large datasets. At the same time, the resulting learned index of the logarithmic PGM, $h_{\varepsilon}^{\prime}$, can be considerably more complex on geometrically structured data.
From learned indices to indexing.
Finally, we considered the following question:
“Can a dynamic learned index be converted into a dynamic indexing data structure?”
To this end, we designed a hybrid technique combining hashing-based fast-access data structures with a doubly linked list to support indexing queries. Our method offers output-sensitive worst-case guarantees, even in the presence of deletions. As it is known that traditional indexing structures currently outperform learned indices in the general dynamic setting [36], we focused our comparisons on improving the theoretical and practical performance within the class of learned approaches.
In practice, the contiguous memory access of the logarithmic PGM index offsets its overhead from querying multiple structures, making it faster in all scenarios, except for an adversarial one. This means that the lower complexity of our learned index does not immediately translate to improved efficiency in the indexing data structure. This raises an interesting open question of whether memory-access efficient fully dynamic approaches to convert a learned index into a dynamic indexing structure can exist.
While we acknowledge that our update-times are slow in comparison with state-of-the art, our approach does illustrate that it brings worst-case guarantees: as it has an advantage when the query-to-update ratio is large and the index has undergone sufficiently many deletions. In adversarial workloads with frequent deletions followed by range queries, we have seen our structure outperform the logarithmic approach – highlighting the value of worst-case guarantees even in specialised settings.
Closing thoughts.
We showed what we believe is an interesting connection between the geometric learned index by Ferragina and Vinciguerra, and dynamic convex hulls from computational geometry. We subsequently provided an implementation of a dynamic learned index that relies on the state-of-the-art dynamic convex hull maintenance algorithms. Our empirical analysis shows that the complex tree rebalancing that is used to dynamically maintain a convex hull currently brings considerable operational overhead compared to low-memory techniques under the logarithmic method. Our experiments, though not uniformly favourable, offer interesting insights into the current barriers and adversarial tradeoffs between worst-case dynamic algorithms and memory-efficient amortised rebuilding schemes.
1:if  $\alpha=$ null OR $\beta=$ null then
2:     return No
3:end if
4:$s(\alpha,\beta)=line(\alpha)\cap line(\beta)$
5:if If $s\in\alpha$ and $s\in\beta$ then
6:     return Yes
7:end if
8:if $\alpha$.slope $<$ $\beta$.slope then
9:     if  $\alpha$.first.$x$ > $s(\alpha,\beta)$.x then
10:         return intersection_test($\alpha$.left, $\beta$)
11:     else if  $\beta$.first.$x$ > $s(\alpha,\beta)$.x  then
12:         return intersection_test($\alpha$, $\beta$.left)
14:         return intersection_test($\alpha$.left, $\beta$)
16:         return intersection_test($\alpha$, $\beta$.left)
17:     else
18:         return yes
19:     end if
20:end if
21:if $\alpha$.slope > $\beta$.slope then
22:     if $\alpha$.second.$x$ $<$ $s(\alpha,\beta).x$ then
23:         return intersection_test($\alpha$.right, $\beta$)
24:     else if $\beta$.second.$x$ $<$ $s(\alpha,\beta).x$ then
25:         return intersection_test($\alpha$, $\beta$.right)
27:         return intersection_test($\alpha$, $\beta$.right)
29:         return intersection_test($\alpha$.right, $\beta$)
30:     else
31:         return yes
32:     end if
33:end if